For the following reaction, calculate how many moles of NO2 forms when 0.356 moles of the...

  1. For the following reaction, calculate how many moles of NO2 forms when 0.356 moles of the reactant completely reacts.

2 N2O5g→4 NO2g+O2g

  1. For the following reaction, calculate how many moles of each product are formed when 0.356 moles of PbS completely react. Assume there is an excess of oxygen.

2PbSs+3O2g→2PbOs+2SO2g

  1. For the following reaction, calculate how many grams of each product are formed when 4.05 g of water is used.

2 H2O→2 H2+O2

4. Determine the theoretical yield of P2O5, when 3.07 g of P reacts with 6.09 g of oxygen in the following chemical equation  

4 P+5O2→2P2O5

  1. Determine the percent yield of the following reaction when 2.80 g of P reacts with excess oxygen. The actual yield of this reaction is determined to by 3.89 g of P2O5.

4 P+5O2→2P2O5

Instructions: Answer the following questions using the following information: ∆Hfus=6.02 kJ/mol; ∆Hvap= 40.7 kJ/mol; specific heat of water is 4.184 J/g∙˚C; specific heat of ice is 2.06 J/g∙˚C; specific heat of water vapor is 2.03 J/g∙˚C.

  1. How much heat is required to vaporize 25 g of water at 100˚C?
  2. How much heat is required to convert 25 g of ice at -4.0 ˚C to water vapor at 105 ˚C (report your answer to three significant figures)?
  3. An ice cube at 0.00 ˚C with a mass of 8.32 g is placed into 55 g of water, initially at 25 ˚C. If no heat is lost to the surroundings, what is the final temperature of the entire water sample after all the ice is melted (report your answer to three significant figures)?

Acid/Base: The Water Constant and pH

Instructions: Determine if each solution is acidic, basic, or neutral.

1.[H3O+] = 1 x 10-10 M; [OH-] = 1 x 10-4 M

  1. [H3O+] = 1 x 10-7 M; [OH-] = 1 x 10-7 M
  2. [H3O+] = 1 x 10-1 M; [OH-] = 1 x 10-13 M
  3. [H3O+] = 1 x 10-13 M; [OH-] = 1 x 10-1 M

Instructions: Calculate [OH-] given [H3O+] in each aqueous solution and classify the solution as acidic or basic.

5. [H3O+] = 2.6 x 10-3 M

6. [H3O+] = 2.6 x 10-8 M

  1. [H3O+] = 1.6 x 10-2 M
  2. [H3O+] = 4.3 x 10-3 M

Instructions: Calculate the [H3O+] of solutions a and b; calculate the [OH-] solutions c and d.

9. pH = 2.76

10. pH = 3.65

  1. pOH = 3.65
  2. pOH = 6.87   

Equilibrium and Le Chateleir's Principal

Instructions: Write the equilibrium expression for each chemical equation.

  1. 2 H2Sg2H2g+S2g
  2. NH4HSsNH3g+H2Sg

Oxidation-Reduction Reactions

Instructions: Assign an oxidation state to each element, ion, or molecule.

  1. Cl-   
  2. Br2
  3. K  
  4. Ca2+

Instructions: Assign an oxidation state to each atom in each compound.

  1. NaCl
  2. CH4
  3. SbCl5
  4. Al2O3

Instructions: Assign an oxidation state to each atom in each polyatomic ion.

  1. NH4+
  2. Cl04-
  3. Cr2072-  
  4. Mn04-

Instructions: Identify the oxidation state of P in each ion.

  1. PO43-   
  2. HPO42-
  3. H2PO4-
  4. H3PO4  
Answer #1

(1) 2N₂Os UNO2 + O2 - 2 mol N2O5 gives u mol NO2 0.356 mol Nels_will give = 4 x 0.356 -10.712 mold 2 (2) 2 Pb Sco) + 302 G 222 P₂O5 . (5) up + 50₂ Mass: 2.8g MW_319 mol initial Moles = 0.09 moles final = 142 goolt - 0.09_x2 0.045. moles = 142 groll mshfus (8.32g) ice - → water water 0°C Dºc - Water (556) 25°C 2 water ToC 1 q2 = q + Dt fus Sesgx 4.184 Ilgºcx (T-25) = 8.329© Com-) = 10-14 - 2.326_x10-12 M. 4.3x10-3 - 3 2_10_2-762_1738_x10-3M ph - (10+) (120+) = -107 10-3-65 = D 2.239 x10-4M- ① ID